# textwidth=0.8textheight=0.9

### v1

@@ 5.3.1 Accuracy Comparison
A major challenge for achieving high accuracy in both PIKANs and PINNs arises when these models are applied to large or wide spatial domains, where numerical instability often occur. To address these challenges, researchers have proposed several strategies, including domain decomposition into overlapping subdomains [54] and regularization techniques such as dropout Tikhonov ensemble Kalman inversion [223]. A particularly effective and conceptually simple alternative is the application of scaling strategies to both the input domain and the PDE residual, mapping them onto a range more compatible with the basis functions, such as the natural interval [1,1][-1,1][ - 1 , 1 ] for Chebyshev polynomials. Mostajeran & Faroughi [68] implemented such a method by scaling both the input coordinates and residual terms in the loss function, resulting in the Scaled-cPIKAN model. This type of scaling helps normalize gradients and stabilize the optimization process, which is especially beneficial in wide spatial domains. They applied this approach to solve the Allen–Cahn equation over the extended domain [6,6][-6,6][ - 6 , 6 ], which is significantly more challenging than the [1,1][-1,1][ - 1 , 1 ] interval. As shown in Table 5, under identical training conditions, Scaled-cPIKAN achieved an 88% reduction in relative error compared to Scaled-PINN and a 90% improvement over the vanilla cPIKAN. Fig. 10(a) illustrates the performance of Scaled-cPIKAN and Scaled-PINN in predicting the Allen–Cahn solution. These results demonstrate that domain and residual scaling not only improves learning efficiency but also enables high accuracy over large spatial domains without increasing model complexity or relying on auxiliary techniques.
@@
Incorporating structured basis functions and architectural enhancements has significantly improved the accuracy of PIKANs. Among the most common basis functions, B-splines have traditionally been used due to their local support and smoothness properties. However, other studies show that structured basis functions such as Chebyshev polynomials [104, 68], wavelets [103, 173], and higher-order ReLU [222] can provide significantly better approximation accuracy than B-splines. Beyond these classical bases, more advanced hybrid kernel-based architectures have been proposed to enhance model accuracy. For example, in [117], Ku̇rková-based KANs (KKANs) replace the MLP layers with structured basis expansions, combining the nonlinear expressiveness of neural networks with the robustness of linear approximation techniques. This design leads to substantial improvements in accuracy without requiring additional techniques such as attention mechanisms. As shown in Table 5, KKAN achieves approximately 67% greater accuracy than the standard PINN, while using only about 8% fewer parameters. Likewise, the Chebyshev-based PIKAN model improves accuracy by roughly 74% compared to PINN, with a comparable parameter count. KKAN demonstrates an accuracy improvement of over 99% compared to the B-spline-based PIKAN, albeit with a parameter count nearly three times higher. These findings highlight the limitations of B-spline bases in complex domains and emphasize the advantages of using structured expansions like Chebyshev polynomials and hybrid architectures such as KKAN. By combining structured linear components with globally expressive basis functions, these models achieve significantly higher accuracy than traditional PINNs, often with comparable or fewer parameters. These results demonstrate the critical role of selecting appropriate basis functions and model architectures in developing accurate and efficient physics-informed neural networks.
@@
Attention mechanisms like Residual-Based Attention (RBA) and its self-scaled variant (ssRBA) can also improve the accuracy of physics-informed models when applied appropriately. RBA works by adaptively weighting regions of the domain using the exponentially weighted moving average of point-wise residuals [231], allowing the model to focus more on areas with larger errors. This attention-like effect enhances accuracy with little computational overhead [55, 152, 232]. The self-scaled version, ssRBA, further improves training stability by dynamically adjusting its internal scaling to preserve a high signal-to-noise ratio during learning [233]. Shukla et al. [66] evaluated RBA in both PINN and PIKAN models. Their findings show that switching from a B-spline to a Chebyshev basis in PIKAN reduces error by over 90% (from $5.83\times 10^{-1}$ to $5.15\times 10^{-2}$) without changing the number of parameters. However, adding RBA to this configuration did not lead to further improvement, indicating that Chebyshev-based PIKAN already achieves high accuracy without requiring attention mechanisms. On the other hand, the PINN model required RBA to achieve a similar level of accuracy, but this improvement came at the cost of using significantly more parameters. The RBA-enhanced PINN used 50049 parameters, which is approximately 645% more than the 6720 parameters used in the Chebyshev-based PIKAN. Fig. 10(b) compares the accuracy of PIKAN and PINN, along with their RBA-invariant counterparts, in predicting the solution of the Allen–Cahn equation over the domain $x\in[-1,1]$ and $t\in[0,1]$. In more advanced architectures like KKAN and Chebyshev-based PIKAN [117], incorporating ssRBA led to very high accuracy gains, up to 99% and 90% error reduction, respectively, but these benefits were achieved by increasing the model size considerably. These results show that PIKAN, especially when using Chebyshev basis functions, can achieve high accuracy with fewer parameters and without relying on attention mechanisms. In comparison, traditional PINNs depend more heavily on RBA or ssRBA to reach similar accuracy, which often requires much larger models.
@@ 5.3.2 Convergence Analysis
Figure 11: Convergence behavior of PIKAN and PINN for solving the Helmholtz equation. (a) Performance comparison of PINN, PIKAN, and cPIKAN variants with and without RBA, trained using L-BFGS and ADAM optimizers on the domain $[-1,1]^{2}$. This panel is adopted from [66]. (b) Training loss evolution of scaled-cPIKAN compared to PINN and unscaled models across different frequency settings on the extended domain $[-4,4]^{2}$. This panel is adopted from [68].
@@
To evaluate convergence behavior, we focus on the loss decay rate, which reflects how quickly the training error decreases over iterations. A faster decay indicates more efficient learning and a more accurate approximation of the solution. Fig. 11(a), adapted from [66], shows the relative $\mathcal{L}^{2}$ error plotted over training iterations and compares the convergence behavior of PIKAN and PINN under both BFGS and Adam optimizers in both parameter-based and computation-based analyses. While PIKAN typically converges more slowly than vanilla PINN, the integration of the RBA strategy substantially improves convergence speed in both architectures. Notably, PINN+RBA and cPIKAN+RBA exhibit the steepest error decay, highlighting the effectiveness of adaptive weighting in accelerating the training process. Fig. 11(b), also adapted from [68], presents loss curves over training iterations to analyze the effects of scaling and recurrence. Scaled-cPIKAN demonstrates the most substantial gains, reducing the loss much faster than its unscaled counterparts. As the wave number increases and the problem complexity grows, Scaled-cPIKAN consistently maintains its superior decay rate, significantly outperforming both vanilla PINN and cPIKAN.
@@ 5.3.3 Spectral Behavior
Figure 12: Comparison of spectral behavior of PIKAN and PINN during training. (a) Evolution of the NTK eigenvalue spectra for different models. cPIKAN shows stable and converging spectra, reflecting well-conditioned training, while PINN and bPIKAN exhibit delayed or unstable spectral dynamics linked to slower or less reliable convergence. (b) Spectral evolution of cPIKAN with varying numbers of temporal subdomains. More subdomains lead to faster spectral convergence and improved conditioning, indicating more efficient and stable training. (c) Mean spectral entropy comparison between cPIKAN and PINN using 8 temporal subdomains. cPIKAN yields consistently lower spectral entropy, suggesting reduced kernel complexity and improved representation of solution features. All panels are adopted from [68]
While the previous sections provided a comparative analysis of PIKAN and PINN regarding accuracy and convergence behavior, this section focuses on a different but equally important aspect of model performance: the spectral characteristics of their training dynamics. Understanding these spectral properties is crucial for explaining the observed convergence trends and for improving the theoretical foundations of physics-informed models. The Neural Tangent Kernel [212, 213, 214] is a particularly valuable tool in this context. Prior work has highlighted the sensitivity of NTK dynamics to network structure and training regimes, especially in the context of PINNs [234, 235]. However, these analyses often rely solely on eigenvalue distributions, which may not capture the full complexity of NTK behavior in KAN models. To address this limitation, Faroughi & Mostajeran [105] have introduced more informative spectral metrics, spectral entropy, a concept derived from Shannon’s information theory [236]. Spectral entropy quantifies the uncertainty or dispersion in the NTK eigenvalue spectrum and is defined as,
@@
where $\lambda_{i}$ are the eigenvalues of the NTK matrix. This metric serves as a proxy for how learning signals are distributed, offering insights into the alignment between model structure and the frequency content of the solution. Fig. 12 summarizes the spectral behavior of the NTK matrix under various settings, complementing empirical observations and providing a theoretical foundation for understanding when and why PIKAN outperforms traditional PINNs.
@@
Analyzing the eigenvalue spectrum of the NTK matrices during training provides critical insight into the learning dynamics and numerical stability of physics-informed based methods. In foundational work, the NTK of PINNs was derived under the infinite-width assumption [234], revealing that the kernel converges to a nearly constant form during training. However, the eigenvalue distribution showed a pronounced spectral bias, causing imbalanced convergence across different components of the loss. This limitation was further explored in [235], where PINNs were shown to struggle with learning initial and boundary conditions in advection- or diffusion-dominated regimes, prompting the use of adaptive loss weighting and periodic activations. More recently, NTK analysis has been extended to Kolmogorov–Arnold-type networks. A simplified study [68] examined the effect of Chebyshev bases on the NTK spectrum using non-nested approximations, while a more comprehensive approach in [105] derived the full NTK for cPIKANs, incorporating nested structures and physics-informed constraints to reflect realistic training setups. Fig. 12(a) presents the evolution of NTK eigenvalues during training for the Helmholtz equation, under a setup that includes domain and residual scaling. The cPIKAN model exhibits stable and smoothly converging eigenvalue spectra across training iterations, indicating consistent kernel behavior and well-conditioned learning dynamics. In contrast, the bPIKAN model displays highly irregular and disordered spectra, with no clear convergence pattern throughout training. This chaotic spectral behavior supports previous findings that bPIKAN fails to accurately learn the Helmholtz solution under the same configuration. For the PINN model, the NTK eigenvalues remain unstable during early training and only begin to stabilize after several thousand iterations. Convergence in the spectrum becomes apparent mainly in the later stages of training, suggesting delayed and uneven learning across different modes. These observations underscore the importance of both the chosen basis and architectural design in shaping NTK dynamics and the general learning behavior during training, particularly for stiff or oscillatory PDEs such as the Helmholtz equation.
@@
An important strategy for improving the spectral behavior and training performance of PINN and PIKAN models is domain decomposition, a technique that partitions the input space, e.g., time or space-time domain, into smaller subdomains to simplify the learning task. Previous studies have highlighted the impact of this approach on spectral bias and convergence dynamics. In particular, Saadat et al. [235] showed that restricting the training domain to a smaller subregion significantly improved the performance of PINNs. Their findings confirmed that spectral bias, caused by the imbalance in learning different solution modes, can be mitigated by reducing the complexity of the target function. This principle underlies several sequential learning strategies, such as time-marching or sequence-to-sequence PINNs, where the model is trained incrementally over smaller time windows. However, such methods come with added algorithmic and computational overhead. More recently, Faroughi & Mostajeran [105] conducted a detailed NTK analysis of cPIKAN models under temporal domain decomposition. Their results demonstrated that as the number of subdomains increases, the NTK matrices become better conditioned and converge more rapidly, leading to more stable learning and improved accuracy. This effect was especially pronounced in diffusion-dominated problems like the Allen–Cahn equation. As illustrated in Fig. 12(b), splitting the time domain into 1, 2, or 4 subdomains results in progressively smoother NTK eigenvalue spectra, confirming the benefits of decomposition in reducing spectral complexity. These findings suggest that domain decomposition is a powerful yet computationally intensive tool for addressing spectral challenges in physics-informed models.
@@
Given the increasing complexity of NTK matrices in advanced physics-informed models, there is a growing need for such advanced spectral metrics that go beyond conventional eigenvalue analysis. Traditional approaches may fail to capture key aspects of training behavior, especially in models involving higher-order derivatives or multi-component loss functions. Wang et al. [234] analyzed the failure of standard PINNs in solving a 1D wave equation and found a severe imbalance in the convergence rates of different loss components, leading to instability during training. By tracking the dynamic evolution of NTK-weighted loss terms, they demonstrated that customized optimization strategies could restore balance and ensure stable learning. In a more challenging case, Faroughi & Mostajeran [105] studied the forced vibration equation, which includes higher-order terms such as $u_{xxxx}$ and $u_{tt}$, along with complex initial and boundary conditions. Instead of relying solely on NTK eigenvalue trajectories, they used spectral entropy, Eq. (20), to assess how concentrated or dispersed the NTK spectrum is during training. As shown in Fig. 12(c), both PINN and cPIKAN are applied to this problem using 8 temporal subdomains. The figure reports the mean spectral entropy of NTK blocks like $K_{uu}$ (solution interactions) and $K_{u_{t}u_{t}}$ (time derivative interactions), showing that cPIKAN consistently achieves lower entropy, indicating better alignment with high-frequency solution features. These results underscore the value of spectral entropy as a diagnostic tool for revealing the deeper dynamics of learning in complex physics-informed models.
6 Deep Operator KANs

### v2

@@ 5.3.1 Accuracy Comparison
A major challenge for achieving high accuracy in both PIKANs and PINNs arises when these models are applied to large or wide spatial domains, where numerical instability often occur. To address these challenges, researchers have proposed several strategies, including domain decomposition into overlapping subdomains howard2024finite and regularization techniques such as dropout Tikhonov ensemble Kalman inversion gao2025scalable . A particularly effective and conceptually simple alternative is the application of scaling strategies to both the input domain and the PDE residual, mapping them onto a range more compatible with the basis functions, such as the natural interval [1,1][-1,1] for Chebyshev polynomials. Mostajeran & Faroughi mostajeran2025scaled implemented such a method by scaling both the input coordinates and residual terms in the loss function, resulting in the Scaled-cPIKAN model. This type of scaling helps normalize gradients and stabilize the optimization process, which is especially beneficial in wide spatial domains. They applied this approach to solve the Allen–Cahn equation over the extended domain [6,6][-6,6], which is significantly more challenging than the [1,1][-1,1] interval. As shown in Table 5, under identical training conditions, Scaled-cPIKAN achieved an 88% reduction in relative error compared to Scaled-PINN and a 90% improvement over the vanilla cPIKAN. Fig. 10(a) illustrates the performance of Scaled-cPIKAN and Scaled-PINN in predicting the Allen–Cahn solution. These results demonstrate that domain and residual scaling not only improves learning efficiency but also enables high accuracy over large spatial domains without increasing model complexity or relying on auxiliary techniques.
@@
Incorporating structured basis functions and architectural enhancements has significantly improved the accuracy of PIKANs. Among the most common basis functions, B-splines have traditionally been used due to their local support and smoothness properties. However, other studies show that structured basis functions such as Chebyshev polynomials guo2024physics , mostajeran2025scaled , wavelets patra2024physics , mostajeran2023novel , and higher-order ReLU so2024higher can provide significantly better approximation accuracy than B-splines. Beyond these classical bases, more advanced hybrid kernel-based architectures have been proposed to enhance model accuracy. For example, in toscano2025kkans , Ku̇rková-based KANs (KKANs) replace the MLP layers with structured basis expansions, combining the nonlinear expressiveness of neural networks with the robustness of linear approximation techniques. This design leads to substantial improvements in accuracy without requiring additional techniques such as attention mechanisms. As shown in Table 5, KKAN achieves approximately 67% greater accuracy than the standard PINN, while using only about 8% fewer parameters. Likewise, the Chebyshev-based PIKAN model improves accuracy by roughly 74% compared to PINN, with a comparable parameter count. KKAN demonstrates an accuracy improvement of over 99% compared to the B-spline-based PIKAN, albeit with a parameter count nearly three times higher. These findings highlight the limitations of B-spline bases in complex domains and emphasize the advantages of using structured expansions like Chebyshev polynomials and hybrid architectures such as KKAN. By combining structured linear components with globally expressive basis functions, these models achieve significantly higher accuracy than traditional PINNs, often with comparable or fewer parameters. These results demonstrate the critical role of selecting appropriate basis functions and model architectures in developing accurate and efficient physics-informed neural networks.
@@
Attention mechanisms like Residual-Based Attention (RBA) and its self-scaled variant (ssRBA) can also improve the accuracy of physics-informed models when applied appropriately. RBA works by adaptively weighting regions of the domain using the exponentially weighted moving average of point-wise residuals anagnostopoulos2024residual , allowing the model to focus more on areas with larger errors. This attention-like effect enhances accuracy with little computational overhead toscano2024inferring , rigas2024adaptive , chen2024self . The self-scaled version, ssRBA, further improves training stability by dynamically adjusting its internal scaling to preserve a high signal-to-noise ratio during learning mcclenny2023self . Shukla et al. shukla2024comprehensive evaluated RBA in both PINN and PIKAN models. Their findings show that switching from a B-spline to a Chebyshev basis in PIKAN reduces error by over 90% (from $5.83\times 10^{-1}$ to $5.15\times 10^{-2}$) without changing the number of parameters. However, adding RBA to this configuration did not lead to further improvement, indicating that Chebyshev-based PIKAN already achieves high accuracy without requiring attention mechanisms. On the other hand, the PINN model required RBA to achieve a similar level of accuracy, but this improvement came at the cost of using significantly more parameters. The RBA-enhanced PINN used 50049 parameters, which is approximately 645% more than the 6720 parameters used in the Chebyshev-based PIKAN. Fig. 10(b) compares the accuracy of PIKAN and PINN, along with their RBA-invariant counterparts, in predicting the solution of the Allen–Cahn equation over the domain $x\in[-1,1]$ and $t\in[0,1]$. In more advanced architectures like KKAN and Chebyshev-based PIKAN toscano2025kkans , incorporating ssRBA led to very high accuracy gains, up to 99% and 90% error reduction, respectively, but these benefits were achieved by increasing the model size considerably. These results show that PIKAN, especially when using Chebyshev basis functions, can achieve high accuracy with fewer parameters and without relying on attention mechanisms. In comparison, traditional PINNs depend more heavily on RBA or ssRBA to reach similar accuracy, which often requires much larger models.
Figure 11: Influence of hyperparameter selection on the relative error of the PIKAN model for the Allen–Cahn example. (a) Layer and neuron configuration. The minimum relative error $(10^{-2})$ occurs in the fifth layer with 10 neurons. (b) Degree and layer configuration. The minimum relative error $(10^{-2})$ is achieved in the fifth layer with a polynomial degree of 5. (c) Number of parameters and residual points configuration. The minimum relative error $(10^{-2})$ is obtained with 2,580 parameters and 50,000 residual points. All panels present new analysis based on mostajeran2025scaled .
Hyperparameter configuration plays a key role in determining both the accuracy and generalization performance of KAN-based models. Despite its importance, only a few studies have systematically examined how these factors influence learning behavior, including zeng2024kan , pourkamali2024kolmogorov . For example, pourkamali2024kolmogorov presented an empirical investigation within a data-driven framework, analyzing the sensitivity of KANs to key hyperparameters such as network depth and spline order. Their findings revealed that, unlike MLPs whose accuracy remained relatively stable with increasing depth, KAN performance tended to degrade as additional layers were introduced. Moreover, the polynomial order of the B-spline activation functions was identified as a crucial factor: higher spline orders generally led to reduced accuracy across real-world datasets. To further illustrate the practical implications of hyperparameter sensitivity, a quantitative analysis is provided for the Allen–Cahn example using the cPIKAN model. Fig. 11 provides a standardized illustration of how variations in key hyperparameters, such as network configuration and the number of residual points, affect the relative $\mathcal{L}^{2}$ error in the PIKAN model. This example demonstrates that accuracy is sensitive to architectural and sampling choices, highlighting the needs for incorporating standardized datasets, metrics, and tuning procedures across studies to ensure comparability, reproducibility, and meaningful benchmarking within the KAN research community.
@@ 5.3.2 Convergence Analysis
Figure 12: Convergence behavior of PIKAN and PINN for solving the Helmholtz equation. (a) Performance comparison of PINN, PIKAN, and cPIKAN variants with and without RBA, trained using L-BFGS and ADAM optimizers on the domain $1,1]^{2}$. This panel is adopted from shukla2024comprehensive . (b) Training loss evolution of scaled-cPIKAN compared to PINN and unscaled models across different frequency settings on the extended domain $4,4]^{2}$. This panel is adopted from mostajeran2025scaled .
@@
To evaluate convergence behavior, we focus on the loss decay rate, which reflects how quickly the training error decreases over iterations. A faster decay indicates more efficient learning and a more accurate approximation of the solution. Fig. 12(a), adapted from shukla2024comprehensive , shows the relative $\mathcal{L}^{2}$ error plotted over training iterations and compares the convergence behavior of PIKAN and PINN under both BFGS and Adam optimizers in both parameter-based and computation-based analyses. While PIKAN typically converges more slowly than vanilla PINN, the integration of the RBA strategy substantially improves convergence speed in both architectures. Notably, PINN+RBA and cPIKAN+RBA exhibit the steepest error decay, highlighting the effectiveness of adaptive weighting in accelerating the training process. Fig. 12(b), also adapted from mostajeran2025scaled , presents loss curves over training iterations to analyze the effects of scaling and recurrence. Scaled-cPIKAN demonstrates the most substantial gains, reducing the loss much faster than its unscaled counterparts. As the wave number increases and the problem complexity grows, Scaled-cPIKAN consistently maintains its superior decay rate, significantly outperforming both vanilla PINN and cPIKAN.
@@ 5.3.3 Spectral Behavior
Figure 13: Comparison of spectral behavior of PIKAN and PINN during training. (a) Evolution of the NTK eigenvalue spectra for different models. cPIKAN shows stable and converging spectra, reflecting well-conditioned training, while PINN and bPIKAN exhibit delayed or unstable spectral dynamics linked to slower or less reliable convergence. (b) Spectral evolution of cPIKAN with varying numbers of temporal subdomains. More subdomains lead to faster spectral convergence and improved conditioning, indicating more efficient and stable training. (c) Mean spectral entropy comparison between cPIKAN and PINN using 8 temporal subdomains. cPIKAN yields consistently lower spectral entropy, suggesting reduced kernel complexity and improved representation of solution features. All panels are adopted from mostajeran2025scaled .
While the previous sections provided a comparative analysis of PIKAN and PINN regarding accuracy and convergence behavior, this section focuses on a different but equally important aspect of model performance: the spectral characteristics of their training dynamics. Understanding these spectral properties is crucial for explaining the observed convergence trends and for improving the theoretical foundations of physics-informed models. The Neural Tangent Kernel jacot2018neural , seleznova2022analyzing , seleznova2022neural is a particularly valuable tool in this context. Prior work has highlighted the sensitivity of NTK dynamics to network structure and training regimes, especially in the context of PINNs Wang2020NTKPinns , saadat2022neural . However, these analyses often rely solely on eigenvalue distributions, which may not capture the full complexity of NTK behavior in KAN models. To address this limitation, Faroughi & Mostajeran faroughi2025neural have introduced more informative spectral metrics, spectral entropy, a concept derived from Shannon’s information theory shannon1948mathematical . Spectral entropy quantifies the uncertainty or dispersion in the NTK eigenvalue spectrum and is defined as,
@@
where $\lambda_{i}$ are the eigenvalues of the NTK matrix. This metric serves as a proxy for how learning signals are distributed, offering insights into the alignment between model structure and the frequency content of the solution. Fig. 13 summarizes the spectral behavior of the NTK matrix under various settings, complementing empirical observations and providing a theoretical foundation for understanding when and why PIKAN outperforms traditional PINNs. The architectures operate in the finite-width regime, as the parameter initialization follows standard scaling, and the NTK evolves during training, capturing feature-learning effects characteristic of practical networks.
@@
Analyzing the eigenvalue spectrum of the NTK matrices during training provides critical insight into the learning dynamics and numerical stability of physics-informed based methods. In foundational work, the NTK of PINNs was derived under the infinite-width assumption Wang2020NTKPinns , revealing that the kernel converges to a nearly constant form during training. However, the eigenvalue distribution showed a pronounced spectral bias, causing imbalanced convergence across different components of the loss. This limitation was further explored in saadat2022neural , where PINNs were shown to struggle with learning initial and boundary conditions in advection- or diffusion-dominated regimes, prompting the use of adaptive loss weighting and periodic activations. More recently, NTK analysis has been extended to Kolmogorov–Arnold-type networks. A simplified study mostajeran2025scaled examined the effect of Chebyshev bases on the NTK spectrum using non-nested approximations, while a more comprehensive approach in faroughi2025neural derived the full NTK for cPIKANs, incorporating nested structures and physics-informed constraints to reflect realistic training setups. Fig. 13(a) presents the evolution of NTK eigenvalues during training for the Helmholtz equation, under a setup that includes domain and residual scaling. The cPIKAN model exhibits stable and smoothly converging eigenvalue spectra across training iterations, indicating consistent kernel behavior and well-conditioned learning dynamics. In contrast, the bPIKAN model displays highly irregular and disordered spectra, with no clear convergence pattern throughout training. This chaotic spectral behavior supports previous findings that bPIKAN fails to accurately learn the Helmholtz solution under the same configuration. For the PINN model, the NTK eigenvalues remain unstable during early training and only begin to stabilize after several thousand iterations. Convergence in the spectrum becomes apparent mainly in the later stages of training, suggesting delayed and uneven learning across different modes. These observations underscore the importance of both the chosen basis and architectural design in shaping NTK dynamics and the general learning behavior during training, particularly for stiff or oscillatory PDEs such as the Helmholtz equation.
@@
An important strategy for improving the spectral behavior and training performance of PINN and PIKAN models is domain decomposition, a technique that partitions the input space, e.g., time or space-time domain, into smaller subdomains to simplify the learning task. Previous studies have highlighted the impact of this approach on spectral bias and convergence dynamics. In particular, Saadat et al. saadat2022neural showed that restricting the training domain to a smaller subregion significantly improved the performance of PINNs. Their findings confirmed that spectral bias, caused by the imbalance in learning different solution modes, can be mitigated by reducing the complexity of the target function. This principle underlies several sequential learning strategies, such as time-marching or sequence-to-sequence PINNs, where the model is trained incrementally over smaller time windows. However, such methods come with added algorithmic and computational overhead. More recently, Faroughi & Mostajeran faroughi2025neural conducted a detailed NTK analysis of cPIKAN models under temporal domain decomposition. Their results demonstrated that as the number of subdomains increases, the NTK matrices become better conditioned and converge more rapidly, leading to more stable learning and improved accuracy. This effect was especially pronounced in diffusion-dominated problems like the Allen–Cahn equation. As illustrated in Fig. 13(b), splitting the time domain into 1, 2, or 4 subdomains results in progressively smoother NTK eigenvalue spectra, confirming the benefits of decomposition in reducing spectral complexity. These findings suggest that domain decomposition is a powerful yet computationally intensive tool for addressing spectral challenges in physics-informed models.
@@
Given the increasing complexity of NTK matrices in advanced physics-informed models, there is a growing need for such advanced spectral metrics that go beyond conventional eigenvalue analysis. Traditional approaches may fail to capture key aspects of training behavior, especially in models involving higher-order derivatives or multi-component loss functions. Wang et al. Wang2020NTKPinns analyzed the failure of standard PINNs in solving a 1D wave equation and found a severe imbalance in the convergence rates of different loss components, leading to instability during training. By tracking the dynamic evolution of NTK-weighted loss terms, they demonstrated that customized optimization strategies could restore balance and ensure stable learning. In a more challenging case, Faroughi & Mostajeran faroughi2025neural studied the forced vibration equation, which includes higher-order terms such as $u_{xxxx}$ and $u_{tt}$, along with complex initial and boundary conditions. Instead of relying solely on NTK eigenvalue trajectories, they used spectral entropy, Eq. (20), to assess how concentrated or dispersed the NTK spectrum is during training. As shown in Fig. 13(c), both PINN and cPIKAN are applied to this problem using 8 temporal subdomains. The figure reports the mean spectral entropy of NTK blocks like $K_{uu}$ (solution interactions) and $K_{u_{t}u_{t}}$ (time derivative interactions), showing that cPIKAN consistently achieves lower entropy, indicating better alignment with high-frequency solution features. These results underscore the value of spectral entropy as a diagnostic tool for revealing the deeper dynamics of learning in complex physics-informed models.
6 Deep-operator KANs